Hi! Please help me with this age related problem: A professor is three times as old as one of his pupil, Freddie. Four years ago he was four times as old as Freddie was at that time. How old is Freddie? Thank you so much!
Answer by stanbon(75887)

A professor is three times as old as one of his pupil, Freddie. Four years ago he was four times as old as Freddie was at that time. How old is Freddie?
------
Equations:
p = 3f
p-4 = 4(f-4)
------------------
Substitute for "p" and solve for "f"::
3f-4 = 4f-16
f = 12 (Freddie age now)
----------------
p = 3f = 36 (Professor's age now)
-----------------
Cheers,
Stan H.
----------------